OpenStudy (anonymous):

Use the product rule to find the derivative of the following y=(3x^2+2)(2x-1)

OpenStudy (vivek3461):

If y = f(x).g(x) then dy/dx = f(x).g'(x) + f '(x).g(x) So for this question, dy/dx = (3x^2+2).2 + (6x)(2x-1)

OpenStudy (anonymous):

so will it be (9x+2)x(2-1)+(9x+2)(2-1)

OpenStudy (vivek3461):

Answer is dy/dx = (3x^2+2).2 + (6x)(2x-1)

OpenStudy (anonymous):

how did you got that?

OpenStudy (vivek3461):

Assume f(x)=3x^2+2, so f '(x)=6x + 0 and g(x)=2x-1, so g'(x)= 2 - 0

OpenStudy (vivek3461):

Use the product rule for final answer

OpenStudy (anonymous):

how did you got 6x+0

OpenStudy (vivek3461):

derivative of X^n is n.X^(n-1) and derivative of a constant is zero